OpenStudy (noah690):

Please help with the following: 1.Write the equation y=-1/2x + 3 in stand form. 2. Write the equation 9x + 2y = 6 in slope-intercept form. 3. Write the equation y-3=4(x+1) in standard form. 4.Write the equation y+7=-2(x-3) in slope-intercept form. 5. Write the equation y-5=3(x-2) is written in point-slope form. Please provide in A) slope-intercept form and B) standard form. Thanks for the help.

OpenStudy (texaschic101):

slope intercept form : y = mx + b standard form : Ax + By = C -- it cannot contain fractions y = -1/2x + 3 --- when you have a fraction, and you do not want one, multiply the entire equation by the common denominator, which in this case, is 2. 2(y = -1/2x + 3) = 2y = -1x + 6 -- add x to both sides x + 2y = 6 -- standard form 9x + 2y = 6 -- we want y term by itself, so we subtract 9x from both sides 9x - 9x + 2y = -9x + 6 -- combine terms 2y = -9x + 6 -- divide both sides by 2 y = -9/2x + 6/2 -- reduce if possible y = -9/2x + 3 Now lets see what you can do with the rest of them.

OpenStudy (noah690):

Thanks for the help Texaschic101.